a grindstone spinning at the rate of 8.3 rev/s has what approximate angular speed

A grindstone spinning at a rate of 8.3 rev/s has what approximate angular speed
Answer #1

1)

As the velocity remains constant and the motion is circular, so -

V1.3 rad sec 0.02

So, option (c) is correct.

----------------------------------------------------

2)

\omega = 8.3 \: rev/sec = \frac{8.3 \times 2\pi\:\: rad}{sec} = 52.15 \:\: rad/sec

So, option (c) is correct.

----------------------------------------------------

3)

Angular acceleration is -

\alpha = \frac{\omega_{f} - \omega_{i}}{\Delta t} = \frac{11-2}{5.5} = 1.6364 \:\: rad/sec^2

So, option (c) is correct.

----------------------------------------------------